OpenStudy (anonymous):

Simplify the expression. Xsqrt(8x^3)+Xsqrt(2)+4sqrt(2x) If the simplified expression is written in standard form, what is the leading coefficient?

OpenStudy (pfenn1):

Is this the expression?\[x \sqrt{8x^3}+x \sqrt2+4\sqrt{2x}\]

OpenStudy (anonymous):

yes

OpenStudy (pfenn1):

Can you simplify \[\sqrt{8x^3}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

type your question using this website. Helps me a lot http://www.myalgebra.com/algebra_solver.aspx